### v1

@@ 10.2 Synthesizing Incompleteness
Figure 11: Examples of tree height map collected in Poznan. (The examples are color mapped and resized for clear visualization. The height increases as the color becomes more yellow.)
@@ 11.3 Sampling Step Analysis
Figure 12: Example of reconstructing corrupted height map with line strip pattern scan using RoofDiffusion.
Figure 13: Samples in PoznanRD dataset.
@@ 11.5 Roof Variety
Figure 13 illustrates the variety of roof types in the PoznanRD dataset. Constructions of valid roofs of different types and with a variety of features have been shown for RoofDiffusion and No-FP RoofDiffusion: gables (Fig. 18(e)), dormers (Figs. 16(a) and 16(h)), shads (Figs. 16(e) and 18(e)), leans (Figs. 15(b) and 15(i)), and flats (Fig. 18(a)).
@@ 11.7 Failure Cases
Figure 14 provides some examples of failure cases where RoofDiffusion misclassifies tree noise as roof structures. LABEL:fig:fail-a and LABEL:fig:fail-b illustrate erroneous reconstructions of unlikely slope and flat planes, respectively, caused by extensive occlusion from tree points. LABEL:fig:fail-c incorrectly identifies tree points as thin wall structures. LABEL:fig:fail-d and LABEL:fig:fail-e mistakenly interpret tree noise as non-existent chimney structures.
@@ 11.9 Additional Qualitative Experimental Results
Figure 17: Evaluation of completion and denoising on the Dales3D dataset [82].
@@
Figure 18: Evaluation of completion and denoising on USGS 3DEP LiDAR data sampled over Wayne County, MI [78].
@@
Figure 19: Evaluation of completion and denoising on USGS 3DEP LiDAR data sampled over Cambridge, MA [79].
@@
Figure 20: Evaluation of completion and denoising for No-FP RoofDiffusion on USGS 3DEP LiDAR data sampled over Wayne County, MI [78]. Compared with linear, pNCNN [15], and CU-Net [85].
@@
Figure 21: Evaluation of completion and denoising for No-FP RoofDiffusionon USGS 3DEP LiDAR data sampled over Cambridge, MA [79].

### v2

@@ 9.2 Synthesizing Incompleteness
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 11: Examples of tree height map in PoznanRDdataset. (The examples are color mapped and resized for clear visualization. The height increases as the color becomes more yellow.)
@@ 10.3 Sampling Step Analysis
Figure 12: RoofDiffusion reconstruction of height maps corrupted by scan line strip patterns.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 13: Samples in the PoznanRD (Poznan Roof Dataset).
@@ 10.5 Roof Variety
Figure 13 illustrates the variety of roof types in the PoznanRD dataset. Constructions of valid roofs of different types and with a variety of features have been shown for RoofDiffusion and No-FP RoofDiffusion: gables (Fig. 19(e)), hips (Figs. 17(a) and 17(h)), shads (Figs. 17(e) and 19(d)), leans (Figs. 16(b) and 16(i)), and flats (Fig. 19(a)).
@@ 10.7 Failure Cases
Figure 14 provides some examples of failure cases where RoofDiffusion misclassifies tree noise as roof structures. Figures 14(a) and 14(b) illustrate erroneous reconstructions of unlikely slope and flat planes, respectively, caused by extensive occlusion from tree points. Figure 14(c) incorrectly identifies tree points as thin wall structures. Figures 14(d) and 14(e) mistakenly interpret tree noise as non-existent chimney structures.
@@ 10.9 Additional Qualitative Experimental Results
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption (a) Refer to caption (b) Refer to caption (c) Refer to caption (d) Refer to caption (e) Refer to caption (f) Refer to caption (g) Refer to caption (h)
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption (a) Refer to caption (b) Refer to caption (c) Refer to caption (d) Refer to caption (e) Refer to caption (f) Refer to caption (g) Refer to caption (h) Refer to caption (i)
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption (a) Refer to caption (b) Refer to caption (c) Refer to caption (d) Refer to caption (e) Refer to caption (f) Refer to caption (g) Refer to caption (h) Refer to caption (i)
Figure 17: Evaluation of the completion and denoising on the Dales3D dataset [82].
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption (a) Refer to caption (b) Refer to caption (c) Refer to caption (d) Refer to caption (e) Refer to caption (f) Refer to caption (g) Refer to caption (h) Refer to caption (i)
Figure 18: Evaluation of the completion and denoising on USGS 3DEP LiDAR data sampled over Wayne County, MI [78].
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption (a) Refer to caption (b) Refer to caption (c) Refer to caption (d) Refer to caption (e) Refer to caption (f) Refer to caption (g) Refer to caption (h) Refer to caption (i)
Figure 19: Evaluation of the completion and denoising on USGS 3DEP LiDAR data sampled over Cambridge, MA [79].
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption (a) Refer to caption (b) Refer to caption (c) Refer to caption (d) Refer to caption (e) Refer to caption (f) Refer to caption (g) Refer to caption (h) Refer to caption (i)
Figure 20: Evaluation of the completion and denoising for No-FP RoofDiffusion on USGS 3DEP LiDAR data sampled over Wayne County, MI [78]. Comparison with linear, pNCNN [15], and CU-Net [85].
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption (a) Refer to caption (b) Refer to caption (c) Refer to caption (d) Refer to caption (e) Refer to caption (f) Refer to caption (g) Refer to caption (h) Refer to caption (i)
Figure 21: Evaluation of the completion and denoising for No-FP RoofDiffusionon USGS 3DEP LiDAR data sampled over Cambridge, MA [79].